# AeroVerse: UAV-Agent Benchmark Suite for Simulating, Pre-training, Finetuning, and Evaluating Aerospace Embodied World Models

### v1

@@ VII-A Quantitative Analysis
3d-llm[9] 0.108 0.0312 0.2808 0.0851 0.0081 0.2171 0.1263 0.0 0.292 0.0609 0.0425 0.2135
gpt-4-vision-preview[20] 0.1277 0.1871 0.3263 0.1193 0.1514 0.3643 0.049 0.0529 0.3320 0.0718 0.0836 0.3392
gpt-4o[20] 0.2343 0.2861 0.4741 0.2137 0.2004 0.3714 0.1349 0.1289 0.5960 0.1473 0.1684 0.4519
blip2-flan-t5-xxl[35] 0.080 0.0611 0.383 0.0686 0.0671 0.3357 0.0788 0.064 0.4100 0.1024 0.0433 0.2789
blip2-opt-6.7b[35] 0.0245 0.0228 0.1400 0.0092 0.004 0.1333 0.0221 0.018 0.1939 0.0358 0.0335 0.1789
instructblip-flan-t5-xxl[38] 0.0345 0.0248 0.3430 0.0324 0.0427 0.2845 0.0207 0.0367 0.4720 0.0637 0.0213 0.3462
instructblip-vicuna-7b[38] 0.1844 0.2168 0.2350 0.1812 0.1576 0.2231 0.1187 0.09 0.1674 0.1364 0.155 0.2000
instructblip-vicuna-13b[38] 0.1917 0.2146 0.2855 0.1952 0.1874 0.3175 0.1245 0.1028 0.2800 0.1508 0.1963 0.2765
llama-adapter-v2-7B[39] 0.2324 0.2692 0.3829 0.2067 0.1976 0.3000 0.1366 0.1191 0.4348 0.1789 0.194 0.3217
llava-v1.5-vicuna-7b[32] 0.1868 0.229 0.3688 0.1461 0.1575 0.2786 0.0727 0.0762 0.4660 0.1297 0.1274 0.3500
llava-v1.5-vicuna-13b[32] 0.1809 0.218 0.3309 0.1701 0.1821 0.3500 0.0971 0.099 0.4490 0.142 0.1532 0.3173
llava-v1.6-vicuna-7b[32] 0.1372 0.1977 0.3242 0.1171 0.1397 0.3571 0.0786 0.0936 0.3939 0.075 0.0977 0.2962
llava-v1.6-vicuna-13b[32] 0.1251 0.1743 0.3503 0.1099 0.1439 0.2952 0.0607 0.060 0.4020 0.0761 0.1002 0.2865
minigpt4[34] 0.2199 0.2969 0.4080 0.1896 0.1978 0.3310 0.1188 0.120 0.5660 0.1374 0.1805 0.3750
mplug[40] 0.1971 0.2485 0.3882 0.1712 0.1729 0.2833 0.0773 0.0842 0.3960 0.0963 0.1248 0.3000
mplug2[41] 0.2127 0.2081 0.4121 0.1881 0.1411 0.3643 0.1125 0.0935 0.4900 0.1474 0.1349 0.3404
qwen-lv-7b[42] 0.1772 0.2152 0.3505 0.1941 0.1837 0.3857 0.1267 0.1278 0.5225 0.1496 0.1128 0.3827
@@
3d-llm[9] 0.0002 0.0019 0.1440 0.0014 0.0027 0.1000 0.0023 0.0021 0.1400 0.0 0.0 0.1044
gpt-4-vision-preview[20] 0.1073 0.0457 0.4520 0.1269 0.0497 0.4067 0.1231 0.0426 0.3711 0.1168 0.0445 0.3867
gpt-4o[20] 0.1064 0.0512 0.5290 0.1152 0.0520 0.5077 0.1167 0.0503 0.4750 0.1183 0.0483 0.5178
blip2-flan-t5-xxl[35] 0.0001 0.0566 0.2000 0.0001 0.0408 0.1844 0.0016 0.0537 0.2200 0.0024 0.0482 0.1956
blip2-opt-6.7b[35] 0.0 0.0002 0.1000 0.0 0.0 0.1000 0.0 0.0 0.1000 0.0 0.0026 0.1044
instructblip-flan-t5-xxl[38] 0.0004 0.0579 0.2356 0.0003 0.0423 0.1875 0.0095 0.0458 0.2733 0.0014 0.072 0.2667
instructblip-vicuna-7b[38] 0.0015 0.0068 0.1386 0.0004 0.0323 0.1546 0.0074 0.0121 0.1667 0.0018 0.0186 0.1625
instructblip-vicuna-13b[38] 0.0012 0.0114 0.1610 0.0034 0.0218 0.1579 0.0063 0.0046 0.1452 0.0007 0.0151 0.1667
llama-adapter-v2-7B[39] 0.0349 0.0216 0.1905 0.0003 0.0 0.1500 0.1138 0.0381 0.2444 0.0838 0.0195 0.1800
llava-v1.5-vicuna-7b[32] 0.1324 0.0339 0.3212 0.1654 0.0357 0.3289 0.1662 0.0376 0.3267 0.1603 0.0414 0.3111
llava-v1.5-vicuna-13b[32] 0.1375 0.0355 0.3640 0.1612 0.038 0.3200 0.1622 0.0448 0.3600 0.1489 0.0497 0.3022
llava-v1.6-vicuna-7b[32] 0.1420 0.0406 03590 0.1697 0.044 0.044 0.1442 0.0382 0.3727 0.1618 0.042 0.3533
llava-v1.6-vicuna-13b[32] 0.1246 0.0399 0.3250 0.1388 0.0391 0.3089 0.1378 0.034 0.3067 0.1535 0.0452 0.3400
minigpt4[34] 0.1288 0.0403 0.2770 0.1396 0.0451 0.2511 0.1689 0.0425 0.2511 0.1585 0.0451 0.2622
mplug[40] 0.1368 0.0368 0.3020 0.1661 0.0397 0.3136 0.1607 0.0393 0.3400 0.1786 0.0409 0.3244
mplug2[41] 0.1426 0.0343 0.3230 0.1687 0.0392 0.3378 0.1674 0.0379 03114 0.1673 0.0442 0.2978
qwen-lv-7b[42] 0.1574 0.0541 0.2850 0.1536 0.0506 0.2956 0.1218 0.0310 0.2800 0.1001 0.0375 0.2178
@@
As presented in Tables I, II, III, and IV, we summarize the overall performance of visual-language models across four UAV downstream tasks within the AeroVerse benchmark. Despite significant advancements in both 2D and 3D visual-language models (VLMs) in recent years, these models continue to encounter challenges with UAV-embodied tasks, including the GPT-4 series. Among the four tasks, existing visual-language models achieve relatively high scores only on SkyAgent-Scene3k, while their performance on the other tasks declines markedly. Overall, gpt-4-vision-review and gpt-4o consistently outperform other models. We will subsequently provide a detailed analysis of the various embodied tasks.
Results on SkyAgent-Scene3k. In evaluating this task, we utilize BLEU, SPICE, and LLM-JUDGE-SCENE to assess the model’s performance in terms of vocabulary richness, semantic accuracy, and human preference. The Qwen-lv-7b model [42] demonstrates strong performance on BLEU, leading in each urban scene, indicating its closer alignment with the reference vocabulary. Overall, the model achieving the highest score in SPICE is gpt-4o, which highlights its advantages in semantic matching. According to the results from LLM-JUDGE-SCENE, the outputs from gpt-4-vision-review and gpt-4o show greater consistency with human preferences.
@@
Results on SkyAgent-Reason3k. For evaluating, we utilize LLM-JUDGE-REASON to assess human preferences. Three models emerge as prominent in this context, i.e., two open-source models, llama-adapter-v2-7B [39] and qwen-lv-7b [42], along with one closed-source model, gpt-4o. A horizontal comparison among the gpt-4 series reveals that gpt-4o demonstrates superior capabilities in first-person spatial reasoning and question-answering tasks.
@@
Results on SkyAgent-Nav3k. Similar to the evaluation metric with SkyAgent-Reason3k, we employ LLM-JUDGE-NAV to assess human preferences. In this task, gpt-4o demonstrate particularly strong performance, ranking at the top across most urban scenarios and evaluation metrics, including vocabulary level, semantic level, and human preference. In the residential area scenario, the response from llama-adapter-v2-7B [39] exhibit greater alignment with the correct answer at the lexical level compared to other models.
@@
Results on SkyAgent-Plan3k. In evaluating this task, we utilize LLM-JUDGE-PLAN to assess human preferences. The performance of many models in this task is notably poor, with several indicators yielding a score of 00. This deficiency arises from the necessity of acquiring a comprehensive range of environmental characteristics for planning long-distance paths. The 3D-LLM struggled to address our inquiries due to its limited generalization capabilities. Although we provide initial-view maps of the true path to the 2D-LLM as environmental information, either through image captions or multi-image input, this information remains incomplete. Notably, gpt-4o ranked first in assessing human preferences across all cities.
VII-B Qualitative Analysis

### v2

@@ VIII-A Quantitative Analysis
3d-llm[Hong20233DLLMIT] 0.108 0.0312 0.2808 0.0851 0.0081 0.2171 0.1263 0.0 0.292 0.0609 0.0425 0.2135
GPT-4-vision-preview[Achiam2023GPT4TR] 0.1277 0.1871 0.3263 0.1193 0.1514 0.3643 0.049 0.0529 0.3320 0.0718 0.0836 0.3392
GPT-4o[Achiam2023GPT4TR] 0.2343 0.2861 0.4741 0.2137 0.2004 0.3714 0.1349 0.1289 0.5960 0.1473 0.1684 0.4519
blip2-flan-t5-xxl[Li2023BLIP2BL] 0.080 0.0611 0.383 0.0686 0.0671 0.3357 0.0788 0.064 0.4100 0.1024 0.0433 0.2789
blip2-opt-6.7b[Li2023BLIP2BL] 0.0245 0.0228 0.1400 0.0092 0.004 0.1333 0.0221 0.018 0.1939 0.0358 0.0335 0.1789
instructblip-flan-t5-xxl[Dai2023InstructBLIPTG] 0.0345 0.0248 0.3430 0.0324 0.0427 0.2845 0.0207 0.0367 0.4720 0.0637 0.0213 0.3462
instructblip-vicuna-7b[Dai2023InstructBLIPTG] 0.1844 0.2168 0.2350 0.1812 0.1576 0.2231 0.1187 0.09 0.1674 0.1364 0.155 0.2000
instructblip-vicuna-13b[Dai2023InstructBLIPTG] 0.1917 0.2146 0.2855 0.1952 0.1874 0.3175 0.1245 0.1028 0.2800 0.1508 0.1963 0.2765
llama-adapter-v2-7B[Gao2023LLaMAAdapterVP] 0.2324 0.2692 0.3829 0.2067 0.1976 0.3000 0.1366 0.1191 0.4348 0.1789 0.194 0.3217
llava-v1.5-vicuna-7b[Liu2023ImprovedBW] 0.1868 0.229 0.3688 0.1461 0.1575 0.2786 0.0727 0.0762 0.4660 0.1297 0.1274 0.3500
llava-v1.5-vicuna-13b[Liu2023ImprovedBW] 0.1809 0.218 0.3309 0.1701 0.1821 0.3500 0.0971 0.099 0.4490 0.142 0.1532 0.3173
llava-v1.6-vicuna-7b[Liu2023ImprovedBW] 0.1372 0.1977 0.3242 0.1171 0.1397 0.3571 0.0786 0.0936 0.3939 0.075 0.0977 0.2962
llava-v1.6-vicuna-13b[Liu2023ImprovedBW] 0.1251 0.1743 0.3503 0.1099 0.1439 0.2952 0.0607 0.060 0.4020 0.0761 0.1002 0.2865
miniGPT4[Zhu2023MiniGPT4EV] 0.2199 0.2969 0.4080 0.1896 0.1978 0.3310 0.1188 0.120 0.5660 0.1374 0.1805 0.3750
mplug[Ye2023mPLUGOwlME] 0.1971 0.2485 0.3882 0.1712 0.1729 0.2833 0.0773 0.0842 0.3960 0.0963 0.1248 0.3000
mplug2[Ye2023mPLUGOwl2RM] 0.2127 0.2081 0.4121 0.1881 0.1411 0.3643 0.1125 0.0935 0.4900 0.1474 0.1349 0.3404
qwen-lv-7b[Bai2023QwenVLAV] 0.1772 0.2152 0.3505 0.1941 0.1837 0.3857 0.1267 0.1278 0.5225 0.1496 0.1128 0.3827
SkyAgent (ours) 0.3660 0.3535 0.4903 0.3519 0.2947 0.4387 0.2962 0.2813 0.5805 0.2961 0.2329 0.3189
SkyAgentX (ours) 0.3669 0.3590 0.4741 0.2976 0.2912 0.4531 0.3466 0.3517 0.5842 0.2253 0.1763 0.3138
@@
3d-llm[Hong20233DLLMIT] 0.0002 0.0019 0.1440 0.0014 0.0027 0.1000 0.0023 0.0021 0.1400 0.0 0.0 0.1044
GPT-4-vision-preview[Achiam2023GPT4TR] 0.1073 0.0457 0.4520 0.1269 0.0497 0.4067 0.1231 0.0426 0.3711 0.1168 0.0445 0.3867
GPT-4o[Achiam2023GPT4TR] 0.1064 0.0512 0.5290 0.1152 0.0520 0.5077 0.1167 0.0503 0.4750 0.1183 0.0483 0.5178
blip2-flan-t5-xxl[Li2023BLIP2BL] 0.0001 0.0566 0.2000 0.0001 0.0408 0.1844 0.0016 0.0537 0.2200 0.0024 0.0482 0.1956
blip2-opt-6.7b[Li2023BLIP2BL] 0.0 0.0002 0.1000 0.0 0.0 0.1000 0.0 0.0 0.1000 0.0 0.0026 0.1044
instructblip-flan-t5-xxl[Dai2023InstructBLIPTG] 0.0004 0.0579 0.2356 0.0003 0.0423 0.1875 0.0095 0.0458 0.2733 0.0014 0.072 0.2667
instructblip-vicuna-7b[Dai2023InstructBLIPTG] 0.0015 0.0068 0.1386 0.0004 0.0323 0.1546 0.0074 0.0121 0.1667 0.0018 0.0186 0.1625
instructblip-vicuna-13b[Dai2023InstructBLIPTG] 0.0012 0.0114 0.1610 0.0034 0.0218 0.1579 0.0063 0.0046 0.1452 0.0007 0.0151 0.1667
llama-adapter-v2-7B[Gao2023LLaMAAdapterVP] 0.0349 0.0216 0.1905 0.0003 0.0 0.1500 0.1138 0.0381 0.2444 0.0838 0.0195 0.1800
llava-v1.5-vicuna-7b[Liu2023ImprovedBW] 0.1324 0.0339 0.3212 0.1654 0.0357 0.3289 0.1662 0.0376 0.3267 0.1603 0.0414 0.3111
llava-v1.5-vicuna-13b[Liu2023ImprovedBW] 0.1375 0.0355 0.3640 0.1612 0.038 0.3200 0.1622 0.0448 0.3600 0.1489 0.0497 0.3022
llava-v1.6-vicuna-7b[Liu2023ImprovedBW] 0.1420 0.0406 03590 0.1697 0.044 0.3756 0.1442 0.0382 0.3727 0.1618 0.042 0.3533
llava-v1.6-vicuna-13b[Liu2023ImprovedBW] 0.1246 0.0399 0.3250 0.1388 0.0391 0.3089 0.1378 0.034 0.3067 0.1535 0.0452 0.3400
miniGPT4[Zhu2023MiniGPT4EV] 0.1288 0.0403 0.2770 0.1396 0.0451 0.2511 0.1689 0.0425 0.2511 0.1585 0.0451 0.2622
mplug[Ye2023mPLUGOwlME] 0.1368 0.0368 0.3020 0.1661 0.0397 0.3136 0.1607 0.0393 0.3400 0.1786 0.0409 0.3244
mplug2[Ye2023mPLUGOwl2RM] 0.1426 0.0343 0.3230 0.1687 0.0392 0.3378 0.1674 0.0379 03114 0.1673 0.0442 0.2978
qwen-lv-7b[Bai2023QwenVLAV] 0.1574 0.0541 0.2850 0.1536 0.0506 0.2956 0.1218 0.0310 0.2800 0.1001 0.0375 0.2178
SkyAgent (ours) 0.4087 0.2653 0.6175 0.3605 0.2404 0.6016 0.3403 0.2257 0.6441 0.4256 0.2429 0.5821
SkyAgentX (ours) 0.4230 0.2557 0.6217 0.3598 0.2704 0.6083 0.3055 0.2327 0.6457 0.3959 0.2753 0.5885
TABLE V: Consolidated Human Evaluation scores across all models, tasks, and scenarios. Scores are on a scale of 0 to 1. Abbreviations for scenarios are as follows: SH (ShangHai), SZ (ShenZhen), CA (Campus), and RE (Residence).
SkyAgent-Scene3k SkyAgent-Reason3k SkyAgent-Nav3k SkyAgent-Plan3k
[…]
3d-llm[Hong20233DLLMIT] 0.12 0.11 0.14 0.13 0.37 0.60 0.23 0.25 0.34 0.29 0.40 0.25 0.10 0.11 0.12 0.10
gpt-4-vision-preview[Achiam2023GPT4TR] 0.41 0.32 0.30 0.40 0.20 0.30 0.16 0.29 0.40 0.39 0.19 0.30 0.36 0.43 0.46 0.38
gpt-4o[Achiam2023GPT4TR] 0.45 0.38 0.35 0.48 0.40 0.59 0.22 0.75 0.70 0.68 0.44 0.57 0.35 0.41 0.45 0.39
blip2-flan-t5-xxl[Li2023BLIP2BL] 0.53 0.46 0.50 0.68 0.19 0.31 0.14 0.28 0.28 0.25 0.30 0.41 0.10 0.11 0.12 0.11
blip2-opt-6.7b[Li2023BLIP2BL] 0.54 0.51 0.49 0.67 0.17 0.20 0.21 0.19 0.14 0.12 0.13 0.18 0.10 0.10 0.10 0.10
instructblip-flan-t5-xxl[Dai2023InstructBLIPTG] 0.58 0.47 0.51 0.66 0.28 0.45 0.18 0.38 0.17 0.17 0.15 0.27 0.11 0.11 0.14 0.11
instructblip-vicuna-7b[Dai2023InstructBLIPTG] 0.61 0.49 0.48 0.72 0.12 0.14 0.20 0.27 0.58 0.59 0.38 0.52 0.12 0.11 0.13 0.11
instructblip-vicuna-13b[Dai2023InstructBLIPTG] 0.60 0.50 0.52 0.71 0.10 0.15 0.11 0.10 0.60 0.63 0.39 0.59 0.11 0.12 0.13 0.10
llama-adapter-v2-7B[Gao2023LLaMAAdapterVP] 0.19 0.17 0.21 0.25 0.41 0.57 0.29 0.74 0.69 0.66 0.45 0.65 0.20 0.11 0.44 0.30
llava-v1.5-vicuna-7b[Liu2023ImprovedBW] 0.18 0.16 0.17 0.22 0.30 0.40 0.19 0.59 0.59 0.49 0.28 0.50 0.40 0.53 0.52 0.47
llava-v1.5-vicuna-13b[Liu2023ImprovedBW] 0.17 0.15 0.18 0.21 0.34 0.41 0.21 0.55 0.57 0.56 0.34 0.55 0.42 0.52 0.54 0.45
llava-v1.6-vicuna-7b[Liu2023ImprovedBW] 0.14 0.13 0.13 0.17 0.18 0.35 0.15 0.34 0.43 0.38 0.29 0.32 0.42 0.57 0.49 0.48
llava-v1.6-vicuna-13b[Liu2023ImprovedBW] 0.15 0.12 0.12 0.16 0.19 0.28 0.15 0.32 0.39 0.36 0.24 0.33 0.38 0.46 0.48 0.46
minigpt4[Zhu2023MiniGPT4EV] 0.21 0.20 0.23 0.23 0.35 0.42 0.17 0.49 0.65 0.61 0.38 0.53 0.39 0.47 0.59 0.47
mplug[Ye2023mPLUGOwlME] 0.16 0.14 0.19 0.19 0.35 0.52 0.27 0.60 0.62 0.57 0.29 0.39 0.41 0.54 0.53 0.53
mplug2[Ye2023mPLUGOwl2RM] 0.20 0.19 0.20 0.28 0.38 0.48 0.26 0.68 0.64 0.60 0.37 0.57 0.43 0.56 0.58 0.50
qwen-lv-7b[Bai2023QwenVLAV] 0.63 0.55 0.52 0.70 0.36 0.49 0.33 0.62 0.55 0.62 0.42 0.58 0.46 0.50 0.46 0.34
SkyAgent (ours) 0.87 0.89 0.88 0.86 0.86 0.84 0.81 0.82 0.84 0.84 0.82 0.81 0.88 0.89 0.87 0.89
SkyAgentX (ours) 0.86 0.85 0.87 0.82 0.88 0.85 0.83 0.89 0.85 0.80 0.86 0.76 0.90 0.88 0.84 0.86
@@
As presented in Tables I, II, III, and IV, we summarize the overall performance of visual-language models across four UAV downstream tasks within the AeroVerse benchmark. Despite significant advancements in both 2D and 3D visual-language models (VLMs) in recent years, these models continue to encounter challenges with UAV-agent embodied tasks, including the GPT-4 series. Among the four tasks, existing visual-language models achieve relatively high scores only on SkyAgent-Scene3k, while their performance on the other tasks declines markedly. Overall, GPT-4-vision-review and GPT-4o consistently outperform other models. However, our SkyAgentX trained by aerospace embodied chain-of-thought and multitask curriculum learning mechanisms has demonstrated breakthrough improvements across multiple metrics. To further validate these results, we also conduct a comprehensive human evaluation, detailed in Table V, which confirms that outputs from our models are overwhelmingly preferred by human judges. We will subsequently provide a detailed analysis of the various embodied tasks.
Results on SkyAgent-Scene3k. The evaluation of lexical richness, semantic accuracy, and human preference is conducted using BLEU, SPICE, and LLM-JUDGE-SCENE metrics. Qwen-lv-7b performes exceptionally well in BLEU, while GPT-4o led in SPICE, demonstrating its advantage in semantic matching. Notably, SkyAgent achieves a leapfrog improvement in both BLEU (Shanghai 0.4302 vs. Qwen 0.2305) and SPICE (0.3083 vs. GPT-4o 0.1114), indicating that the aerospace embodied chain-of-thought mechanism significantly enhances lexical and semantic expression capabilities. However, its LLM-JUDGE-SCENE score remains slightly lower than the GPT-4 series, suggesting that further optimization is needed for human preference alignment. Nevertheless, our direct human evaluation tells a different story. As shown in Table V, SkyAgent achieves a human preference score of 0.87 in the Shanghai scenario, far surpassing GPT-4o (0.45). This indicates that human evaluators strongly prefer the detailed and accurate descriptions generated by our model, a nuance not fully captured by the automated LLM-JUDGE metric.
@@
Results on SkyAgent-Reason3k. For evaluating, we utilize LLM-JUDGE-REASON to assess human preferences. Three models emerge as prominent in this context, i.e., two open-source models, llama-adapter-v2-7B [Gao2023LLaMAAdapterVP] and qwen-lv-7b [Bai2023QwenVLAV], along with one closed-source model, GPT-4o. A horizontal comparison among the GPT-4 series reveals that GPT-4o demonstrates superior capabilities in first-person spatial reasoning and question-answering tasks. SkyAgent achieves BLEU/SPICE scores of 0.4598/0.3846 in the Shanghai scenario, significantly outperforming the baseline model. After introducing aerospace embodied multitask curriculum learning, the REA score improves further, demonstrating that progressive learning enhances the model’s understanding of complex spatial relationships. In the residence scenario, the curriculum learning variant (i.e., SkyAgentX) achieves a BLEU score of 0.4399, a 23% increase over the original version (i.e., SkyAgent), highlighting its adaptability to local environmental characteristics. This quantitative superiority is strongly corroborated by our human evaluation (Table V), where SkyAgentX achieves a score of 0.88 in Shanghai, decisively outperforming all baseline models, including GPT-4o (0.40), proving a significant human preference for our model’s reasoning abilities.
@@
Results on SkyAgent-Nav3k. Similarly to the evaluation metric with SkyAgent-Reason3k, we employ LLM-JUDGE-NAV to assess human preferences. In this task, GPT-4o demonstrates particularly strong performance, ranking at the top across most urban scenarios and evaluation metrics, including vocabulary level, semantic level, and human preference. In the residential area scenario, the response from llama-adapter-v2-7B [Gao2023LLaMAAdapterVP] exhibit greater alignment with the correct answer at the lexical level compared to other models. SkyAgent achieves an NAV score of 0.4903 in the Shanghai navigation task, with its explicit reasoning capability being particularly suitable for problems involving spatial reasoning. Crucially, the human evaluation results in Table V underscore our model’s practical advantage. SkyAgentX scores 0.85 in Shanghai, significantly higher than GPT-4o’s 0.70. This suggests that the navigational instructions generated by our model are clearer, more accurate, and more useful to a human user, even when automated metrics show closer competition.
@@
Results on SkyAgent-Plan3k. In evaluating this task, we utilize LLM-JUDGE-PLAN to assess human preferences. The performance of many models in this task is notably poor, with several indicators yielding a score of 0. This deficiency arises from the necessity of acquiring a comprehensive range of environmental characteristics for planning long-distance paths. The 3D-LLM struggled to address our inquiries due to its limited generalization capabilities. Although we provide initial-view maps of the true path to the 2D-LLM as environmental information, either through image captions or multi-image input, this information remains incomplete. Through the aerospace embodied chain-of-thought mechanism of drone agent, SkyAgent achieves a PLAN score of 0.6175 in Shanghai tests, significantly surpassing GPT-4o. When combined with aerospace embodied multitask curriculum learning strategy, the planning capability of SkyAgentX improves further and delivers the best performance in the tests in four cities. This remarkable performance is overwhelmingly supported by human evaluations (Table V). In the Shanghai scenario, SkyAgentX receives a score of 0.90 from human judges, highlighting its exceptional ability to generate coherent and viable long-range plans that are vastly superior to those from any other model, including GPT-4o (0.35). Thanks to the phased environmental complexity learning mechanism, the model demonstrates significantly enhanced stability in long-path planning for complex scenarios, providing key technical support for autonomous navigation exploration in real-world settings.
VIII-B Qualitative Analysis